Ana is 5 years more than Jack. The sum of their ages is 29. Find the ages of Ana and Jack ?
Answer by lenny460(1073)

Ana is 5 years more than Jack. The sum of their ages is 29. Find the ages of Ana and Jack ?
Let the age of Jack be x years.
Let the age of Ana be (x + 5) years.
Therefore:
x + (x + 5) = 29
2x + 5 = 29
2x = 24
x = 12
x + 5 = 17

Answer:
Ana is 17 years and Jack is 12 years.




lennox obuong
algebra tutor